For 1980 through 1995, the annual sales, S (in thousands of dollars) of a photography studio can be modeled by S = -1/8t^2 + 3t +42, where t=0 represents 1980. During which year did the studio sell $59,500?
Answer by nerdybill(7384)

For 1980 through 1995, the annual sales, S (in thousands of dollars) of a photography studio can be modeled by S = -1/8t^2 + 3t +42, where t=0 represents 1980. During which year did the studio sell $59,500?
.
Set S to 59.5 and solve for t:
59.5 = -1/8t^2 + 3t +42
476 = -t^2 + 24t + 336
0 = -t^2 + 24t - 140
0 = t^2 - 24t + 140
0 = (t-14)(t-10)
t = {10, 14}
.
1980+10 = 1990
and
1980+14 = 1994
.
Ans: 1990 and 1994